OpenStudy (anonymous):

type this equation in standard form. 3x=6y+8 type this equation in standard form. 3x=6y+8 @Mathematics

OpenStudy (anonymous):

y=1/2x-4/3

OpenStudy (anonymous):

err maybe not. is it 3x-6y=8

OpenStudy (anonymous):

3x-6y-8=0

OpenStudy (anonymous):

the standard form is ax+by+c=0 so 3x-6y-8=0

OpenStudy (anonymous):

i have forgotten all of the forms haha.

OpenStudy (anonymous):

someone should look at my question please

OpenStudy (anonymous):

thanks guy

OpenStudy (anonymous):

well cum